OpenStudy (anonymous):

Why is the integral of 1/(2x+1) =.5ln(2x+1) and not just ln(2x+1)

OpenStudy (noelgreco):

Try differentiating ln(2x+1). Don't forget the chain rule. Note the 2 in the numerator. That .5 cancels out the 2.

OpenStudy (anonymous):

hmm ic.

OpenStudy (phi):

or to say it this way: \[\int\limits \frac{1}{u} du = \ln u\] in your case u = 2x+1 and du = 2 dx you need 2 dx so you write it like this \[ \frac{1}{2} \int \frac{1}{2x+1} 2 dx \]

OpenStudy (anonymous):

alrite, thanks.